10. (8 pt) The light intensity from the reference cell in a dual-beam spectrometer (Po) is 0.50 W/cm2. The light intensity from the sample (P) is 0.025 W/cm. a. What is the transmittance of the sample? b. What is the absorbance of the sample? After opening the instrument, the cover is left open at a crack letting some stray light hit the detector at 0.010 W/cm2 c. What is the apparent absorbance as a result of the stray light?
Answer #1

10. a) Tranamittance (T) = I/Io =

I = intensity of the light after passing through sample. I0 = intensity of the incident light.

T = 0.025/0.5 = 0.05

B) absorbance A = - log T = - log 0.05 = 1.30

C) when stray light is used

Incident intensity and intensity after absorption both changes

SL is intensity of stray light.

Absorbance (A) = -log (I + SL)/(I0+ SL)

= -log ( 0.025+0.010/0.5+0.010)

= - log( 0.035/0.51) = 1.163

D)

Perecentage error = |(measured - accepted)| *100/ accepted

( 1.30-1.163)*100/1.30 = 10.53 %